# Generating non-Clifford gate operations through exact mapping between Majorana fermions and $\mathbb{Z}_{4}$ parafermions

### v2

@@ I Introduction
Since the end of the last century, the concept of quantum computing has emerged as a paradigmatic idea which utilizes quantum physics machinery to perform computational tasks Grover1996 ; Shor1994 ; Deutsch1985 ; Feynman1982 . Specifically, rather than storing and processing data as classical bits that consist of ``$0"$s and ``$1"$s, quantum computing involves the definition and manipulation of quantum bits (qubits), which are quantum states living in a $2^{n}$ Hilbert space with $n$ being some positive integer. It has been envisioned that a functional quantum computer is capable of solving tasks that otherwise would take more than a lifetime in under one second Madsen2022 .
Building a full-fledged quantum computer is of no easy feat. Due to the lack of natural error-correction mechanisms in quantum mechanical systems, qubits are prone to errors that are typically ubiquitous in nature due to, e.g., interaction with environment and/or hardware imperfection Kitaev2001 ; Papic2023 ; Riste2014 . While quantum error-correction algorithms exist and have been extensively studied, their success require an already very small error in the system. For this reason, much effort has been devoted towards designing a physical system that hosts qubits with minimal errors. Topological systems are potentially one of the most promising platforms for this purpose due their natural robustness that results from nonlocality. Indeed, the idea of using topological systems for quantum computing has emerged as an active area on its own which is dubbed as topological quantum computing Pachos2017 ; Sarma2015 ; Fre2008 .
@@
Interest in $\mathbb{Z}_{n}$ parafermions has resulted in extensive theoretical studies that revolve around detecting them in experiments and utilizing them for quantum computations Liu2022 ; Kaskela2021 ; Fendley2012 ; Clarke2013 ; Mong2014 . While the proposals for realizing general $\mathbb{Z}_{n}$ parafermions often involve access to the elusive fractional quantum Hall system, restricting ourselves to the specific type of $\mathbb{Z}_{4}$ parafermions opens some alternative, more realistic, approaches Chew2018 ; Calzona2018 ; Orth2015 ; Zhang2014 . Indeed, since a mapping exists between $\mathbb{Z}_{4}$ parafermions and Majorana fermions, see, e.g., Ref. Bomantara2021 , $\mathbb{Z}_{4}$ parafermions could in principle be realized in platforms that support Majorana fermions with the addition of appropriate interaction terms Chew2018 ; Calzona2018 ; Orth2015 ; Zhang2014 and/or periodic driving Bomantara2021 .
@@
This paper is organized as follows. In Sec. II, we describe the mathematical definitions that are used in the treatment of Majorana fermions and their generalization to parafermions (focusing on $\mathbb{Z}_{4}$ parafermions). Specifically, it covers Majorana fermions and parafermions defining operators, parity operators, projectors, measurement operators, braiding operators, as well as their mapping to Pauli matrices and generalized Pauli matrices respectively. Section III, which is further broken down into three parts, focuses on the mapping between a system of $4$ Majorana fermions and a system $2$ $\mathbb{Z}_{4}$ parafermions without enforcing a total parity constraint. Section III.1 mathematically elucidates the exact mapping between the two systems. Section III.2 then exhaustively presents the quantum gate operations that can be achieved by braiding every pair among the $4$ Majorana fermions system from the perspective of Pauli matrices and their $4\times 4$ generalizations. Similarly, Sec. III.3 exhaustively presents the quantum gate operations that can be achieved by braiding the pair of $\mathbb{Z}_{4}$ parafermions from both perspectives. Section IV, which is also broken into three parts, develops the exact mapping between a system of $6$ Majorana fermions and a system of $4$ $\mathbb{Z}_{4}$ parafermions under a conserved total parity. Section III.1 mathematically details the exact mapping between the two systems. Section IV.2 then exhaustively presents the quantum gate operations that can be achieved by braiding every pair among the $6$ Majorana fermions while preserving the total parity from the perspective of Pauli matrices and their $4\times 4$ generalizations. Similarly, Sec. IV.3 exhaustively presents the quantum gate operations that can be achieved by braiding every pair among the $4$ $\mathbb{Z}_{4}$ parafermions from both perspectives. Section V elucidates a measurement based protocol to realize the braiding of Majorana fermions and $\mathbb{Z}_{4}$ parafermions. Finally, Sec. VI summarizes the results and outlines future research directions.
@@ II.1 Overview of Majorana fermions
which is unitary, squares to identity, and consequently has two eigenvalues of ±$\pm 1$ (see Appendix A). A projector on the ±$\pm 1$ eigenstate of the parity can then be written as
@@
It can be written as
@@ II.2 Overview of $\mathbb{Z}_{n}$ parafermions
which is a unitary operator, equal to identity only when raised to the power of $n$, and consequently has a set of eigenvalues equal to $W_{n}$ (see Appendix B). Other definitions presented below specialize to the case of $\mathbb{Z}_{4}$, which is the main subject of this manuscript. The eigenstate associated with the $k\in W_{n}$ eigenvalue of parity operator will be denoted by |$\ket{k}$ in the following. A projector on |$\ket{s}$ can then be written as
@@
[…]
@@
A unitary operator acting on an $N$-dimensional ($N$D) qudit is said to be a Clifford operator if it transforms Pauli gates, under conjugation, into Pauli gates. Mathematically this group of operators can be expressed as Tolar ,
@@
For the special case of $2^{n}$ dimensional qubits, Clifford gates include a multiplicative combination of the $S$ gate, Hadamard gate, and CNOT gate Gottesman . An $N$D generalization of the $S$ and Hadamard gates can be written as, respectively, Tolar
@@ III Quantum gate operations achieved by braiding without parity conservation constraint
For simplicity, we consider the minimal system of either $4$ Majorana fermions or $2$ $\mathbb{Z}_{4}$ parafermions. These $4$ Majorana fermions are labeled as $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$, whilst the equivalent $2$ parafermions are $\psi_{1}$ and $\psi_{2}$. In the absence of parity conservation constraint, the Hilbert space spanned by the Majorana fermions is four dimensional and could be described by the following effective Pauli matrices,
@@ III.2 Braiding Majorana Fermions
There are $6$ possible braidings with $4$ Majorana fermions, they are
@@ III.3 Braiding parafermions
In the minimal system of $2$ $\mathbb{Z}_{4}$ parafermions, the only possible braiding operator is
@@
Unsurprisingly, such a braiding leads only to a diagonal quantum gate operation in the qudit space. Interestingly, when viewed from the two-qubit perspective spanned by $4$ Majorana fermions, the resulting quantum gate operation does not belong to the Clifford gate set for qubits, as it contains the sought-after $T$ gate, i.e.,
$U^{(P)}=-\frac{1+i}{\sqrt{2}}\sigma_{z}^{(2)}\cdot\sigma_{z}^{(1)}\cdot H^{(1)}\cdot CNOT_{2,1}\cdot H^{(1)}\cdot({T^{(1)}})^{\dagger}.$ (31)
Therefore, conversely to the observation of Sec. III.2, braiding of $\mathbb{Z}_{4}$ parafermions may also lead to universal gate operations on qubits spanned by Majorana fermions.
@@ IV Quantum gate operations achieved by braiding under parity conservation constraint
These conservation laws respectively reduce the logical Hilbert space of our system of Majorana fermions or parafermions to $4$ dimensions. Consequently, physical operators involving $6$ Majorana fermions, i.e., parities, projectors, and braiding operators, could be written in terms of 4 $\mathbb{Z}_{4}$ parafermions and vice versa in such a total parity conserving setting.
@@
$\sigma_{z}^{(2)}=i\gamma_{4}\gamma_{5}$ , $\sigma_{x}^{(2)}=i\gamma_{4}\gamma_{6}.$ (34)
It can be easily checked that these operators commute with $P_{\rm tot}^{(MF)}$ and therefore respect the total parity constraint. The same Hilbert space could be equivalently described by the parafermions via the following effective generalized Pauli matrices,
@@ IV.2 Braiding Majorana fermions
There are totally $15$ possible braidings with $6$ Majorana fermions. The $6$ braiding operators from Sec. III.2 now correspond to quantum gate operations
@@
The difference in the resulting quantum gate operation as compared in Sec. III.2 when braiding the same pair of Majorana fermions is attributed to the fact that the effective Pauli matrices here take different forms from those of Sec. III.2 so that they are compatible with the total parity conservation constraint. Nevertheless, it is worth noting that the quantum gate operations arising from these braiding operations also belong to the Clifford set. For completeness, we present the braiding operators for the remaining $8$ pairs of Majorana fermions and their corresponding quantum gate operations in Appendix D.
@@
$U_{13}^{(MF)}\cdot\mathcal{T}_{x}\cdot\left(U_{13}^{\dagger}\right)^{(MF)}=\frac{1}{2}(\mathcal{T}_{x}+\mathcal{T}_{x}^{3}+i\mathcal{T}_{z}^{2}(\mathcal{T}_{x}-\mathcal{T}_{x}^{3})),$ (43)
and is consequently a non-Clifford gate.
@@ IV.3 Braiding Parafermions
All possible braiding operators achievable with $4$ $\mathbb{Z}_{4}$ parafermions reads (See Appendix E for the detailed procedure in obtaining them)
@@
non-Clifford gates can be clearly seen in $U_{1,2}$ that contains a $T$ gate, as well as in
@@ V.1 Measurement-based braiding of Majorana fermions
×$\times(\mathbb{I}+s_{4}\gamma_{3}\gamma_{4})$
@@
Regardless of the measurement outcomes, the result is proportional to the true braiding operator $U_{1}^{(MF)}$ up to products of some Majorana fermions that can be written in terms of appropriate Pauli gates. Therefore, if the Pauli gates are readily accessible in the platform used to enact the series of measurements above, one may simply apply the appropriate Pauli gates depending on the recorded measurement outcomes to retrieve $U_{1}^{(MF)}$ P . Alternatively, due to the symmetry in the measurements involved in the protocol, one may in principle repeat the protocol indefinitely and keep track of all the measurement outcomes until exactly $U_{1}^{(MF)}$ is achieved.
@@ V.2 Measurement-based braiding of parafermions
where $M_{ij}^{(P)}$ is the parafermion parity measurement operator, which is given by Eq. (10) but with the measurement outcome $s$ being randomly selected among $1,i,-1,-i$. Through a similar mechanism as in the Majorana fermions case, this protocol results in either a braiding operator or a braiding operator multiplied by a combination of 4D generalized Pauli matrices depending on the measurement outcomes. Specifically, by leaving the mathematical details in Appendix F, we obtain the following effective unitary operator acting on the main subspace for $s_{1}=s_{4}$,
@@
In particular, we observe that $U_{1}^{(P)}$ (corresponding to the measurement outcomes of $s_{1}=s_{4}$ and $s_{3}=s_{2}$) is already the desired braiding operator. The other unitary operators that arise from obtaining the various different measurement outcomes are related to $U_{1}^{(P)}$ by a product of parafermion operators, which can in turn be expressed in terms of 4D generalized Pauli matrices via the Fradkin-Kadanoff transformation.
VI Concluding remarks
@@
In this paper, the exact mapping between Majorana fermions and $\mathbb{Z}_{4}$ parafermions was developed by focusing on two minimal systems with $4$D Hilbert space. One of these systems consists of either $4$ Majorana fermions or $2$ $\mathbb{Z}_{4}$ parafermions without any constraint on the total parity, whereas the other consists of either $6$ Majorana fermions or $4$ $\mathbb{Z}_{4}$ parafermions under a conserved total parity. The main finding of this paper is the revelation that brading of Majorana fermions may correspond to a non-Clifford quantum gate when viewed from the parafermionic 4D qudit perspective, and vice versa. This result suggests that, in principle, quantum universality could be established in a topological manner in systems of either Majorana fermions or $\mathbb{Z}_{4}$ parafermions if the braiding operations involving one type of quasiparticles (Majorana fermions or $\mathbb{Z}_{4}$ parafermions) are supplemented by braiding of the other type of quasiparticles.
@@
The mapping developed in this work could manifest itself in a physical system comprising one type of quasiparticle through the presence of suitable interaction terms that effectively bind the quasiparticles to form the other type of quasiparticles. As an explicit example, one could consider the system developed in Ref. Bomantara2021 . There, the system describes two chains of Majorana fermions which are subject to appropriately devised periodic driving and interaction terms, so that it supports a pair of $\mathbb{Z}_{4}$ parafermion edge modes. In the absence of the interaction terms, the system would simply reduce to two copies of Kitaev chains Kitaev2001 and instead supports two pairs of Majorana edge modes. In this case, depending on the interaction strength, one could store quantum information either in a $4$D-qudit spanned by two $\mathbb{Z}_{4}$ parafermion edge modes or in a two-qubit state spanned by four Majorana edge modes. Consequently, both braiding of Majorana fermions and parafermion edge modes could, in principle, be executed on demand by tuning the interaction strength and performing the appropriate braiding protocol.
@@
In view of the above, explicitly carrying out the various braiding protocols in the physical system of Ref. Bomantara2021 represents a promising direction for future work. Such braiding protocols could be based on a series of measurements developed in the present work or through an adiabatic means. In the former, each measurement could be carried out by connecting each end of the system to a terminal and measuring the parity-dependent conductance when appropriate voltage bias is applied Bomantara2020 . In the latter, a slowly varying Hamiltonian that enacts braiding could be made by tuning appropriate system parameters such as hopping, on-site potential, interaction strength, and/or superconducting pairing locally. Another interesting aspect for future studies is to devise a system that supports coexisting $Z_{4}$ parafermion and Majorana edge modes. In such a system, richer quantum gate operations are expected due to a potential braiding operation involving a Majorana edge mode and a $Z_{4}$ parafermion, though its precise mathematical definition remains to be established. Finally, it is envisioned that developing a quasiparticle-based quantum chip that could switch its building blocks between Majorana fermions and $Z_{4}$ parafermions on demand would pave the way for harnessing the above obtained non-Clifford gates and subsequently establishing a universal topological quantum computer.

### v3

@@ I Introduction
Since the end of the last century, the concept of quantum computing has emerged as a paradigmatic idea which utilizes quantum physics machinery to perform computational tasks Grover1996 ; Shor1994 ; Deutsch1985 ; Feynman1982 . Specifically, rather than storing and processing data as classical bits that consist of 0``0"s and 1``1"s, quantum computing involves the definition and manipulation of quantum bits (qubits), which are quantum states living in a $2^{n}$-dimensional Hilbert space with $n$ being some positive integer. It has been envisioned that a functional quantum computer is capable of solving tasks that otherwise would take more than a lifetime in under one second Madsen2022 .
Building a full-fledged quantum computer is no easy feat. Due to the lack of natural error-correction mechanisms in quantum mechanical systems, qubits are prone to errors that are typically ubiquitous in nature due to, e.g., interaction with environment and/or hardware imperfection Kitaev2001 ; Papic2023 ; Riste2014 . While quantum error-correction algorithms exist and have been extensively studied, their success require an already very small error in the system. For this reason, much effort has been devoted towards designing a physical system that hosts qubits with minimal errors. Topological systems are potentially one of the most promising platforms for this purpose due their natural robustness that results from nonlocality. Indeed, the idea of using topological systems for quantum computing has emerged as an active area on its own which is dubbed as topological quantum computing Pachos2017 ; Sarma2015 ; Fre2008 .
@@
Interest in $\mathbb{Z}_{n}$ parafermions has resulted in extensive theoretical studies that revolve around detecting them in experiments and utilizing them for quantum computations Liu2022 ; Kaskela2021 ; Fendley2012 ; Clarke2013 ; Mong2014 ; Benhemou2022 . While the proposals for realizing general $\mathbb{Z}_{n}$ parafermions often involve access to the elusive fractional quantum Hall system, restricting ourselves to the specific type of $\mathbb{Z}_{4}$ parafermions opens some alternative, more realistic, approaches Chew2018 ; Calzona2018 ; Orth2015 ; Zhang2014 . Indeed, since a mapping exists between $\mathbb{Z}_{4}$ parafermions and Majorana fermions, see, e.g., Ref. Bomantara2021 , $\mathbb{Z}_{4}$ parafermions could in principle be realized in platforms that support Majorana fermions with the addition of appropriate interaction terms Chew2018 ; Calzona2018 ; Orth2015 ; Zhang2014 and/or periodic driving Bomantara2021 ; Sreejith2016 ; Zhu2022 .
@@
This paper is organized as follows. In Sec. II, we describe the mathematical definitions that are used in the treatment of Majorana fermions and their generalization to parafermions (focusing on $\mathbb{Z}_{4}$ parafermions). Specifically, it covers Majorana fermions and parafermions defining operators, parity operators, projectors, measurement operators, braiding operators, as well as their mapping to Pauli matrices and generalized Pauli matrices respectively. Section III, which is further broken down into three parts, focuses on the mapping between a system of four Majorana fermions and a system two $\mathbb{Z}_{4}$ parafermions without enforcing a total parity constraint. Section III.1 mathematically elucidates the exact mapping between the two systems. Section III.2 then exhaustively presents the quantum gate operations that can be achieved by braiding every pair among the four Majorana fermions system from the perspective of Pauli matrices and their $4\times 4$ generalizations. Similarly, Sec. III.3 exhaustively presents the quantum gate operations that can be achieved by braiding the pair of $\mathbb{Z}_{4}$ parafermions from both perspectives. Section IV, which is also broken into three parts, develops the exact mapping between a system of six Majorana fermions and a system of four $\mathbb{Z}_{4}$ parafermions under a conserved total parity. Section III.1 mathematically details the exact mapping between the two systems. Section IV.2 then exhaustively presents the quantum gate operations that can be achieved by braiding every pair among the six Majorana fermions while preserving the total parity from the perspective of Pauli matrices and their $4\times 4$ generalizations. Similarly, Sec. IV.3 exhaustively presents the quantum gate operations that can be achieved by braiding every pair among the four $\mathbb{Z}_{4}$ parafermions from both perspectives. Section V elucidates a measurement based protocol to realize the braiding of Majorana fermions and $\mathbb{Z}_{4}$ parafermions. Finally, Sec. VI summarizes the results and outlines future research directions.
@@ II.1 Overview of Majorana fermions
which is unitary, squares to identity, and consequently has two eigenvalues of $\pm 1$ (see Appendix A). A projector on the $\pm 1$ eigenstate of the parity can then be written as
@@
It can be written as Ivanov2001 ; Kauffman2004 ; Bravyi2002 ; Padmanabhan2025 :
@@ II.2 Overview of $\mathbb{Z}_{n}$ parafermions
[…]
@@
$\left(U_{jk}^{(P)}\right)^{\dagger}\psi_{j}U_{jk}^{(P)}=-i{\psi_{k}}^{3}{\psi_{j}}^{2}$ $\&$ $\left(U_{jk}^{(P)}\right)^{\dagger}\psi_{k}U_{jk}^{(P)}={\psi_{j}}.$
@@
A unitary operator acting on an $N$-dimensional (NND) qudit is said to be a Clifford operator if it transforms Pauli gates, under conjugation, into Pauli gates. Mathematically this group of operators can be expressed as Tolar ,
@@
For the special case of $2^{n}$ dimensional qubits, Clifford gates include a multiplicative combination of the $S$ gate, Hadamard gate, and CNOT gate Gottesman . An NND generalization of the $S$ and Hadamard gates can be written as, respectively, Tolar
@@ III Quantum gate operations achieved by braiding without parity conservation constraint
For simplicity, we consider the minimal system of either four Majorana fermions or two $\mathbb{Z}_{4}$ parafermions. These four Majorana fermions are labeled as $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, and $\gamma_{4}$, whilst the equivalent two parafermions are $\psi_{1}$ and $\psi_{2}$. In the absence of parity conservation constraint, the Hilbert space spanned by the Majorana fermions is four dimensional and could be described by the following effective Pauli matrices,
@@
Intuitively, $\sigma_{z}^{(1)}$ ($\sigma_{z}^{(2)}$) is related to the occupancy of the number operator associated with a full (complex) fermion formed by $\gamma_{1}$ and $\gamma_{2}$ ($\gamma_{3}$ and $\gamma_{4}$). In this case, the four dimensional Hilbert space could physically represent a system of two complex fermions, one of which is formed by $\gamma_{1}$ and $\gamma_{2}$, whilst the other is formed by $\gamma_{3}$ and $\gamma_{4}$.
@@ III.2 Braiding Majorana Fermions
There are six possible braidings with four Majorana fermions, they are
@@ III.3 Braiding parafermions
In the minimal system of two $\mathbb{Z}_{4}$ parafermions, the only possible braiding operator is
@@
Unsurprisingly, such a braiding leads only to a diagonal quantum gate operation in the qudit space. Interestingly, when viewed from the two-qubit perspective spanned by four Majorana fermions, the resulting quantum gate operation does not belong to the Clifford gate set for qubits, as it contains the sought-after $T$ gate, i.e.,
$U^{(P)}=-\frac{1+i}{\sqrt{2}}\sigma_{z}^{(2)}\cdot\sigma_{z}^{(1)}\cdot H^{(1)}\cdot CNOT_{2,1}\cdot H^{(1)}\cdot({T^{(1)}})^{\dagger},$ (31)
where $T^{(1)}$ is the $T$ gate acting on the first qubit spanned by $\sigma_{x}^{(1)}$ and $\sigma_{z}^{(1)}$. Therefore, conversely to the observation of Sec. III.2, braiding of $\mathbb{Z}_{4}$ parafermions may also lead to universal gate operations on qubits spanned by Majorana fermions.
@@ IV Quantum gate operations achieved by braiding under parity conservation constraint
These conservation laws respectively reduce the logical Hilbert space of our system of Majorana fermions or parafermions to $4$ dimensions. Consequently, physical operators involving six Majorana fermions, i.e., parities, projectors, and braiding operators, could be written in terms of 4 $\mathbb{Z}_{4}$ parafermions and vice versa in such a total parity conserving setting.
@@
$\sigma_{z}^{(2)}=i\gamma_{4}\gamma_{5}$ , $\sigma_{x}^{(2)}=i\gamma_{4}\gamma_{6},$ (34)
which physically represents a system of three complex fermions formed by $\gamma_{1}$ and $\gamma_{2}$, $\gamma_{4}$ and $\gamma_{5}$, and $\gamma_{3}$ and $\gamma_{6}$ respectively, under the conservation of the total number of fermions. It can be easily checked that these operators commute with $P_{\rm tot}^{(MF)}$ and therefore respect the total parity constraint. The same Hilbert space could be equivalently described by the parafermions via the following effective generalized Pauli matrices,
@@ IV.2 Braiding Majorana fermions
There are totally $15$ possible braidings with six Majorana fermions. The six braiding operators from Sec. III.2 now correspond to quantum gate operations
@@
The difference in the resulting quantum gate operation as compared in Sec. III.2 when braiding the same pair of Majorana fermions is attributed to the fact that the effective Pauli matrices here take different forms from those of Sec. III.2 so that they are compatible with the total parity conservation constraint. Nevertheless, it is worth noting that the quantum gate operations arising from these braiding operations also belong to the Clifford set. For completeness, we present the braiding operators for the remaining eight pairs of Majorana fermions and their corresponding quantum gate operations in Appendix D.
@@
$U_{13}^{(MF)}\cdot\mathcal{T}_{x}\cdot\left(U_{13}^{\dagger}\right)^{(MF)}=\frac{1}{2}(\mathcal{T}_{x}+\mathcal{T}_{x}^{3}+i\mathcal{T}_{z}^{2}(\mathcal{T}_{x}-\mathcal{T}_{x}^{3})).$ (43)
Therefore, $U_{13}^{(MF)}$ is a non-Clifford qudit gate.
@@ IV.3 Braiding Parafermions
All possible braiding operators achievable with four $\mathbb{Z}_{4}$ parafermions reads (See Appendix E for the detailed procedure in obtaining them)
@@
non-Clifford gates can be clearly seen in $U_{1,2}$ that contains a $T$ gate ($T^{(1)}$), as well as in
@@ V.1 Measurement-based braiding of Majorana fermions
$\times(\mathbb{I}+s_{4}\gamma_{3}\gamma_{4})$
@@
Regardless of the measurement outcomes, the result is proportional to the true braiding operator $U_{1}^{(MF)}$ up to products of some Majorana fermions that can be written in terms of appropriate Pauli gates. Therefore, if the Pauli gates are readily accessible in the platform used to enact the series of measurements above, one may simply apply the appropriate Pauli gates depending on the recorded measurement outcomes to retrieve $U_{1}^{(MF)}$ P .
Alternatively, due to the symmetry in the measurements involved in the protocol, one may in principle repeat the protocol indefinitely and keep track of all the measurement outcomes until exactly $U_{1}^{(MF)}$ is achieved. Assuming all outcomes of the protocol are equally probable, there is a $1/4$ chance that braiding operation is obtained directly in one cycle of measurements. In general, applying the protocol $23$ times ($92$ measurements in total) yields 99.9% probability (this can be calculated using Binomial Distribution) to achieve the desired braiding. In this case, $92$ is the safe upper bound for the number of measurements needed to realize a braiding operation. In experimental platforms based on InAs/Al heterostructures, recent work Aghaee2024 reports that a single parity measurement could be completed on a scale of a few microseconds. Therefore, the $92$ measurements could be completed in less than a millisecond, the latter of which is a typical quasiparticle poisoning time of the system Aghaee2024 .
@@ V.2 Measurement-based braiding of parafermions
where $M_{ij}^{(P)}$ is the parafermion parity measurement operator, which is given by Eq. (10) but with the measurement outcome $s$ being randomly selected among 1,i,1,i1,i,-1,-i. Through a similar mechanism as in the Majorana fermions case, this protocol results in either a braiding operator or a braiding operator multiplied by a combination of 4D generalized Pauli matrices depending on the measurement outcomes. Specifically, by leaving the mathematical details in Appendix F, we obtain the following effective unitary operator acting on the main subspace for $s_{1}=s_{4}$,
@@
In particular, we observe that $U_{1}^{(P)}$ (corresponding to the measurement outcomes of $s_{1}=s_{4}$ and $s_{3}=s_{2}$) is already the desired braiding operator. The other unitary operators that arise from obtaining the various different measurement outcomes are related to $U_{1}^{(P)}$ by a product of parafermion operators, which can in turn be expressed in terms of 4D generalized Pauli matrices via the Fradkin-Kadanoff transformation. Assuming all outcomes of the protocol are equally probable, applying one cycle of measurements has a $1/16$ chance to yield the desired braiding protocol directly. Applying the protocol $62$ times ($248$ measurements in total) then yields braiding with 98% probability. Assuming that a single measurement of parafermion parity takes around the same time as a single measurement of Majorana fermion parity, $248$ measurements correspond to a timescale that is of the order of hundreds of microseconds in a typical InAs/Al heterostructures Aghaee2024 , which is shorter than the quasiparticle poisoning time of the system, i.e., typically of the order of milliseconds Aghaee2024 .
VI Discussion and concluding remarks
@@
In this paper, the exact mapping between Majorana fermions and $\mathbb{Z}_{4}$ parafermions was developed by focusing on two minimal systems with 44D Hilbert space. One of these systems consists of either four Majorana fermions or two $\mathbb{Z}_{4}$ parafermions without any constraint on the total parity, whereas the other consists of either six Majorana fermions or four $\mathbb{Z}_{4}$ parafermions under a conserved total parity. The main finding of this paper is the revelation that brading of Majorana fermions may correspond to a non-Clifford quantum gate when viewed from the parafermionic 4D qudit perspective, and vice versa. In particular, we have shown explicitly that some braiding operators involving Majorana fermions could be decomposed into 4D qudit gates that include non-Clifford components, such as a single qudit rotation by $\theta=-3\pi/4$. Conversely, we have also explicitly demonstrated the decomposition of some braiding operators involving parafermions in terms of non-Clifford qubit gate components, such as a $T$ gate, a controlled $S$ gate, and/or a controlled rotation gate by $\theta=\pm 3\pi/4$. This result suggests that quantum universality could, in principle, be established in a topological manner in systems of either Majorana fermions or $\mathbb{Z}_{4}$ parafermions if the braiding operations involving one type of quasiparticles (Majorana fermions or $\mathbb{Z}_{4}$ parafermions) are supplemented by braiding of the other type of quasiparticles.
It is worth emphasizing that our main findings above do not contradict with the well-known fact that braiding of Majorana fermions or braiding of parafermions alone does not lead to universality. Indeed, the key observation here is that braiding of parafermions (respectively, Majorana fermions), when expressed in terms of Majorana fermions (respectively, parafermions) through our mapping, cannot be written in terms of braidings of Majorana fermions (respectively, parafermions) only. Given that Majorana fermionic braidings generate $n$-qubit Clifford gates, whereas parafermionic braidings generate $n$-qudit Clifford gates, Majorana fermionic braidings alone or parafermionic braidings alone cannot establish universality, as expected. It is the combination of Majorana fermionic braidings and parafermionic braidings that allows for universality, regardless of which representation (parafermionic or Majorana fermionic) is being used. This is made possible by the fact that one set of braidings only generates Clifford gates, whilst the other set generates some non-Clifford gates. Therefore, both perspectives indeed lead to the same conclusion, i.e., quantum universality, provided both Majorana fermionic braidings and parafermionic braidings could be achieved in the chosen perspective. On the other hand, the presence of only one family of braidings, i.e., Majorana fermionic braidings alone or parafermionic braidings alone, is not sufficient for universality in either representation, as it should be.
@@
The mapping developed in this work could manifest itself in a physical system comprising one type of quasiparticle through the presence of suitable interaction terms that effectively bind the quasiparticles to form the other type of quasiparticles. A periodically driven system is expected to be a promising platform for this purpose. For example, the system developed in Ref. Bomantara2021 describes two chains of Majorana fermions which are subject to appropriately devised periodic driving and Hubbard-like interaction terms, so that it supports a pair of $\mathbb{Z}_{4}$ parafermion edge modes. In the absence of the interaction terms, the system would simply reduce to two copies of Kitaev chains Kitaev2001 and instead supports two pairs of Majorana edge modes. In this case, depending on the interaction strength, one could store quantum information either in a 44D-qudit spanned by two $\mathbb{Z}_{4}$ parafermion edge modes or in a two-qubit state spanned by four Majorana edge modes. Consequently, both braiding of Majorana fermions and parafermion edge modes could, in principle, be executed on demand by tuning the interaction strength and performing the appropriate braiding protocol.
@@
In view of the above, explicitly carrying out the various braiding protocols in the physical system of Ref. Bomantara2021 represents a promising direction for future work. Such braiding protocols could be based on a series of measurements developed in the present work or through an adiabatic means. In the former, each measurement could be carried out by connecting each end of the system to a terminal and measuring the parity-dependent conductance when appropriate voltage bias is applied Bomantara2020 . In the latter, a slowly varying Hamiltonian that enacts braiding could be made by tuning appropriate system parameters such as hopping, onsite potential, interaction strength, and/or superconducting pairing locally.
As precise control over interaction strength might be challenging to achieve in existing experiments, another interesting aspect for future studies is to devise a periodically driven system that supports $Z_{4}$ parafermion and Majorana edge modes simultaneously. In particular, as the coexistence of Majorana edge modes at different quasienergy excitations has been established and utilized for quantum computing applications in recent studies Bomantara2018PRB ; Bomantara2020 ; Bomantara2020s , it is reasonable to expect that such a Majorana-parafermionic hybrid system exists. In such a system, braiding of Majorana and parafermion modes could be potentially carried out under a fixed interaction strength. In this case, the Hilbert space spanned by the Majorana modes and that spanned by the parafermion modes could in principle be coupled in a controllable manner by adapting the adiabatic protocol developed in Ref. Bomantara2018PRB . In particular, this could be achieved by varying the adiabatic parameters every appropriate integer multiple of the driving period, so that the Hilbert spaces spanned by different quasienergy excitations are joined temporarily during the adiabatic process; the Hilbert spaces remain independent outside the application of the protocol, thereby allowing subsequent braiding operations to be carried out without unforeseen complications, whilst quantum information could be transferred from one Hilbert space (e.g., spanned by Majorana fermions) to another (e.g., spanned by parafermions) as needed.
That said, it is worth noting that the aforementioned adiabatic protocol is geometric in nature and not strictly topological, i.e., it may be sensitive to certain types of local noise or parameter fluctuations. Nevertheless, it remains more robust than typical dynamical approaches because it does not require precise control over protocol duration. We also acknowledge that the protocol of Ref. Bomantara2018PRB is applied to the case where the two Hilbert spaces are spanned by Majorana fermions. Therefore, while it should be feasible to extend such a protocol to the case where one Hilbert space is spanned by Majorana fermions and the other by parafermions, the detail of its construction warrants a separate study that is beyond the scope of this present work.
To conclude the above discussion, our findings could in principle be implemented in some periodically driven Majorana-parafermionic hybrid system, the topologically protected braidings of which could be appropriately performed in the Majorana sector or parafermionic sector of the total Hilbert space in a way that does not require applying nonlocal transformations physically. However, the caveat is that to truly establish universality, an additional adiabatic protocol that is geometric (not strictly topological) in nature must be supplemented. Nevertheless, this drawback in our proposed idea is by no means a general drawback to implementing our findings. We remain optimistic that future developments in device design or theory may reveal a fully topological approach for coupling the Majorana fermionic and parafermionic subspaces. In this case, the drawback of our implementation proposal could actually serve as a promising motivation for the further development and studies of a Majorana-parafermionic hybrid system and the means to harness its universality as established in this work. We thus hope that our findings and our implementation proposal may inspire others to explore alternative routes toward this goal.